OpenStudy (anonymous):

Give the equation y = x^2 - 4, what are the solutions to x^2 - 4 = 0 ?

OpenStudy (anonymous):

\[(x+2)(x-2)=0\]

OpenStudy (anonymous):

difference of two squares

OpenStudy (anonymous):

How would you explain that is correct?

OpenStudy (anonymous):

\[a^2-b^2=(a-b)(a+b)\] so\[x=2,x=-2\]